# Uncovering Vulnerabilities of LLM-Assisted Cyber Threat Intelligence

### v2

@@ 4.3 Study of Intertwined Vulnerabilities (RQ6)
Appendix D.2 also details additional analyses with case studies to complement our discussions.
Design Insight . Vulnerabilities studied in this work are not solely caused by model design, but by the inherently fragmented and adversarial threat landscape. Noise such as contradictory proofs-of-concept, reused indicators, or incomplete metadata can propagate into models. Addressing these vulnerabilities therefore requires combining data curation, adversarially aware fine-tuning, and inference-time safeguards to integratively control the adversarial influences.
@@ 5 Related Work
LLM-as-agent. LLM-based agents have been explored across diverse domains such as education (Chu et al., 2025), scientific discovery (Schmidgall et al., 2025), data science (Hong et al., 2024), and urban mobility modeling (Wang et al., 2024). These applications highlight the ability of LLM agents to decompose complex tasks, integrate external tools, and generate executable outputs. In cybersecurity, distributed detection frameworks (Dong et al., 2023) and adaptive rule-evasion defenses (Uetz et al., 2024) illustrate how agent-like systems can strengthen enterprise protection.
LLMs for cybersecurity. LLMs are increasingly leveraged for both offensive and defensive cybersecurity tasks, owing to their strong natural language understanding and reasoning capabilities. In static analysis, LSAST augments traditional SAST tools with dynamic vulnerability knowledge (Keltek et al., 2025), while hybrid systems use LLM-driven preprocessing and explanation to improve anomaly detection in IoT (Ghimire et al., 2025). Autonomous agents have demonstrated the ability to exploit one-day vulnerabilities (Fang et al., 2024b) and guide fuzzing across multi-hop library dependencies (Zhou et al., 2024). Additional applications include enhancing intrusion detection (G. Lira et al., 2024), supporting large-scale code review (Sun et al., 2025), and enabling malware tracking through dataset augmentation and semantic analysis (Yu et al., 2024).
Vulnerabilities of LLMs. Despite their promise, LLM-based agents exhibit critical vulnerabilities that undermine reliability in high-stakes domains. Recent work shows that malfunction amplification attacks can cascade small reasoning errors into severe misjudgments (Zhang et al., 2024). Other studies highlight vulnerabilities to data poisoning (Wang et al., 2025), adversarial prompts Debenedetti et al. (2024), and misalignment (Fang et al., 2024a), all of which raise concerns for deploying LLMs in security-sensitive environments.
@@ 6 Conclusion
This work presents systematic yet intensive studies of intrinsic vulnerabilities that constrain the effectiveness of LLMs in cyber threat intelligence. By combining large-scale benchmark evaluations with real-world CTI reports, we uncover three fundamental failure modes (spurious correlations, contradictory knowledge, and constrained generalization) that persist across multiple CTI stages and workflows. Our autoregressive, human-in-the-loop methodology enables reliable categorization of failure instances and provides insights into how these limitations emerge and propagate. These findings not only reveal blind spots in current LLM reasoning but also chart a path toward more principled model adaptations and robust cyber agent design.
@@ Appendix A Complementary Details of CTI
  • Topic Modeling. Topic modeling techniques such as Latent Dirichlet Allocation (LDA) (Blei et al., 2003), Non-negative Matrix Factorization (Lee & Seung, 1999), and more recent neural topic models (Dieng et al., 2020) are applied to large corpora of unstructured CTI text, including threat reports, incident tickets, and log annotations. These models group documents or paragraphs into coherent themes, enabling analysts to identify clusters of related activity such as “phishing campaigns leveraging Office macros” or “ransomware families exploiting VPN vulnerabilities.” In practice, topic modeling supports early triage by prioritizing threat feeds that share thematic overlap with active campaigns in a given sector.
  • Event Extraction. Event extraction identifies structured incidents—who did what, when, and how—from unstructured logs or reports. Natural Language Processing (NLP) pipelines detect entities such as vulnerabilities (CVEs), indicators of compromise (IP addresses, file hashes), and TTPs (MITRE ATT&CK techniques (Strom et al., 2018)), then associate them with temporal markers and actor actions. In Security Operations Centers (SOCs), event extraction enables long-form reports to be converted into structured JSON objects or STIX bundles (oas, 2017), which can be automatically ingested into SIEM platforms such as Splunk or Elastic. This reduces manual parsing effort and ensures consistent representation across heterogeneous data sources.
  • Knowledge Base Mapping. Knowledge base mapping aligns extracted entities and events with standardized taxonomies such as CVE (CVE Program, 2024), CWE (MITRE Corporation, 2024b), MITRE ATT&CK (Strom et al., 2018), and CAPEC (MITRE Corporation, 2024a). This process typically involves both exact matching (e.g., direct CVE ID resolution) and approximate entity linking (e.g., mapping the phrase “remote PowerShell execution” to ATT&CK T1059.001). In CTI practice, knowledge base mapping is essential for interoperability: intelligence can be shared across organizations using a common language of identifiers, enabling correlation of local incidents with global adversary behaviors.
  • Information Retrieval. Information retrieval systems allow analysts to ground their findings in relevant historical reports and external databases. Implementations include keyword- and embedding-based retrieval from proprietary CTI feeds (e.g., Recorded Future, Mandiant Advantage), open-source repositories (e.g., AlienVault OTX, AbuseIPDB), and structured vulnerability catalogs such as NVD (National Institute of Standards and Technology (2024), NIST) and the CISA KEV catalog (Cybersecurity and Infrastructure Security Agency (2024), CISA). Retrieval pipelines often combine lightweight keyword filters for precision with dense embedding search for semantic coverage. Analysts use these systems to verify whether a newly observed domain has prior associations with known malware campaigns, or whether a vulnerability is actively being exploited.
@@
  • Named Entity Recognition (NER). NER systems are used to extract structured entities such as malware families, infrastructure elements, or actor names from unstructured reports and incident logs. For example, extracting references to APT28, Mimikatz, or QakBot across heterogeneous feeds helps analysts consolidate threat narratives (Barnes & Kordzadeh, 2021).
  • Relation and Event Extraction. Beyond isolated entities, attribution requires uncovering relationships among adversary tactics, techniques, and infrastructure. Relation extraction techniques map, for instance, shared IP ranges between phishing campaigns or code reuse across malware variants, while event extraction captures sequences of actions such as exploitation followed by lateral movement (Husari et al., 2018a; Zhang et al., 2019).
  • Infrastructure Correlation. Adversaries frequently reuse or repurpose command-and-control (C2) infrastructure. CTI teams apply graph-based correlation to link domain registrations, TLS certificates, and hosting providers, enabling attribution of otherwise fragmented observations to known actor toolkits or campaigns (Stone-Gross et al., 2009; Antonakakis et al., 2017).
  • Stylistic and Linguistic Profiling. Analysts also consider linguistic cues in adversary communications or malware code artifacts. Stylometry and compilation fingerprinting methods can identify patterns in variable naming, debugging strings, or grammar usage, which help to tie malware development back to particular groups (Alonso et al., 2019; Rosenblum et al., 2011).
  • Campaign Graph Construction. To synthesize diverse evidence, CTI analysts build structured graphs of adversarial campaigns, linking entities, infrastructure, and TTPs across incidents. Graph construction enables propagation of attribution hypotheses and detection of actor evolution over time (Park et al., 2020; Yang et al., 2021).
@@
  • Historical correlation and trend analysis. Analysts correlate prior incidents and intrusion campaigns to identify recurring attacker playbooks. For example, statistical methods on longitudinal CVE exploitation data help assess whether recently disclosed vulnerabilities follow exploitation trends of past families (Bilge & Dumitras, 2012; Allodi & Massacci, 2014).
  • Exploit prediction scoring. Models such as the Exploit Prediction Scoring System (EPSS) estimate the probability that a vulnerability will be exploited within a given time window, using features such as vulnerability metadata, CVSS scores, and real-world exploit observations (Jacobs & Romanosky, 2019; 2021).
  • Temporal modeling of campaign progression. Recurrent neural networks and temporal point processes capture how campaigns unfold over time, modeling likely transitions from initial access to follow-on payloads such as ransomware (Wang et al., 2019; Okutan & Yilmaz, 2020).
  • Temporal forcasting. LLMs predict which TTPs an actor is likely to employ next or what’s impact. This supports proactive defense, such as generating detection rules for tactics not yet observed in the ongoing campaign (Sood & Enbody, 2013; Husari et al., 2018b).
  • Threat propagation simulation (synthesis). Agent-based and epidemic-style models simulate the spread of malware or worms across interconnected systems, forecasting infection curves and propagation likelihood (Xu et al., 2012; Liu et al., 2015).
@@
  • Mitigation efficacy evaluation. Beyond applying countermeasures, analysts must assess their effectiveness. Approaches include red-teaming exercises, breach-and-attack simulation (BAS) platforms, and adversary emulation scenarios that replay known TTPs to test whether mitigations succeed in preventing, detecting, or containing malicious activity (Shostack, 2014; Peisert et al., 2021).
  • Response playbook recommendation. Structured playbooks standardize incident response by encoding lessons from CTI. These include step-by-step containment and recovery actions tailored to adversary campaigns (e.g., disabling compromised accounts, isolating infected subnets). Orchestration tools such as SOAR platforms automate playbook execution, integrating CTI feeds into dynamic workflows (Mitnick & Vamosi, 2018; Scarfone & Grance, 2012).
  • Summarization. Finally, mitigation intelligence is communicated through concise, contextualized reports for executives and IT operators. Summarization synthesizes prioritized vulnerabilities, mapped mitigations, and recommended workflows. This ensures decision makers understand trade-offs between operational impact and security gains, and supports cross-team coordination in enterprise-scale defense (Mavroeidis & Bromander, 2018; Anwar et al., 2021).

### v3

@@ 4.4.2 Analyzing intertwined vulnerabilities
Appendix D.2 provides additional analyses and case studies to complement this discussion.
5 Potential Defensive Strategy
Based on the experimental insights (§4), we further counteract CTI vulnerabilities at two levels of granularity: (1) a universal intervention applied during inference-time LLM reasoning to defend against all potential vulnerabilities, and (2) a set of fine-grained, qualitative mitigations applied to cyberspace.
5.1 A causal intervention design to validate and mitigate CTI vulnerabilities
Design principle: counterfactual consistency. Despite significant differences, the three vulnerability classes (spurious correlation, contradictory knowledge, and constrained generalization) share a common failure pattern in LLM reasoning: models tend to overfit to surface-level statistical co-occurrences rather than the invariant causal attributes that actually determine the correct CTI judgment.
Definition : Causal attributes refer to CTI factors that should causally drive the label or conclusion, e.g., the attacker’s core TTPs (techniques, procedures, and behavioral signatures), exploitation mechanism, malware or tool behaviors/capabilities, and other stable indicators of the underlying attack process, rather than incidental context such as timestamps, victim industry, geography, or reporting-source artifacts.
To mitigate CTI vulnerabilities, we propose an inference-time counterfactual consistency check ($C^{3}$) based on a simple principle: if a CTI judgment is truly driven by these causal attributes, it should remain stable when we perturb non-causal contextual tokens (e.g., dates, sectors) that should not affect the conclusion. Formally, we model CTI analysis as a mapping $f(x)\rightarrow y$. We then define a set of targeted intervention operators $\mathcal{I}=\{I_{\text{spur}},I_{\text{contra}},I_{\text{gen}}\}$ that generate counterfactual variants $x^{\prime}_{k}$ by editing the non-causal tokens associated with each failure type, while preserving the CTI context (i.e., the causal attributes) that is leveraged to determine $y$.
Validation (diagnosis). To validate the existence of latent vulnerabilities without ground truth labels, we utilize $C^{3}$ as a sensitivity-based detector. For a given input $x$ and a target prediction $y$, we compute a Consistency Score $\Omega$:
$\Omega(y|x)=\frac{1}{|\mathcal{K}|}\sum_{k=1}^{|\mathcal{K}|}\mathbb{1}\left(\text{argmax}_{y^{\prime}}P(y^{\prime}|I_{k}(x))=y\right)$ (1)
where $\mathcal{K}$ is the manifold of generated counterfactuals. If $\Omega(y|x)<\tau$ (a stability threshold), the prediction is flagged as “Vulnerable.” For example, if changing the report’s date causes the model to flip its attribution from an APT identifier to “unknown,” we validate the presence of a contradictory knowledge failure (Type 2.1). This metric provides a quantifiable proxy for model robustness, allowing us to detect specific failure modes in real-time.
Inference-time mitigation (correction). The $C^{3}$ framework goes beyond diagnosis and can serve as an active mitigation engine. Instead of simply relying on the greedy decoding of the original input, we compute the final output $\hat{y}$ via a weighted consensus across the counterfactual set:
$\hat{y}=\text{argmax}_{y}\sum_{k}w_{k}\cdot P(y|x^{\prime}_{k})$ (2)
where $w_{k}$ penalizes paths with high entropy. This effectively filters out reasoning driven by spuriousness (Type 1.x), contradictions (Type 2.x), or overfitting (Type 3.x) by forcing LLM reasoning to converge on evidence that is invariant across contexts.
Experimental results. We empirically show the efficacy of this design by applying the $C^{3}$ intervention across our benchmark tasks. Table 6 details the absolute performance gains ($\Delta$) observed when switching from standard inference to our causal mitigation. We obverse a universal improvement, particularly in tasks sensitive to high-noise metadata. For instance, false flag detection presents a significant improvement (up to +0.115+0.115 F1 for Mixtral), demonstrating that spurious correlation was indeed the primary bottleneck. Similarly, event timeline construction improves significantly across all models (avg $\Delta\approx 0.08$), showing that $C^{3}$ successfully mitigates contradictory knowledge conflicts. Notably, while specialized agentic models show smaller relative gains due to their higher baseline, the intervention still yields consistent improvements, proving that these causal vulnerabilities are inherent to the LLM architecture rather than just domain knowledge gaps.
5.2 Fine-grained mitigation tailored to specific vulnerability subtypes
Building on the causal intervention results above, we now design a set of fine-grained, actionable mitigation strategies tailored to each vulnerability subtype.
5.2.1 Mitigating spurious correlation failures
Spurious correlation failures arise when LLMs overgeneralize from patterns that are statistically associated but not causally linked. These failures typically stem from co-occurrence signals, reused observables, or implicit biases embedded in CTI report structures and metadata. To mitigate such issues, we recommend a combination of (i) retrieval-time evidence filtering and (ii) prompt-level constraints, which incorporate causal reasoning and uncertainty modeling to reduce the reliance on misleading associations and encourage evidence-based reasoning. Below, we detail tailored designs:
1.1 Co-mention bias from raw threat incident. (i) Retrieval: run entity resolution + role labeling, then keep only entities with incident-role evidence (e.g., “operator,” “infrastructure owner,” “victim”) and down-rank “mentioned-in-passing” names. (ii) Inference: require an evidence link for every predicted actor/tool (quote/span + relation type), and add a simple rule: reject if the only support is co-mention.
1.2 Exploitation bias from deliberately reused IoCs. (i) Retrieval: maintain an IoC reuse profile (first/last seen, frequency, and how many different campaigns it appears in) and down-rank “high-reuse” IoCs unless they come with campaign-specific behavioral clues (delivery method, exploit chain, payload traits). (ii) Inference: force a two-step check: first ask “is this IoC discriminative here, or commonly reused?”, then only make attribution/recommendations if it is discriminative.
1.3 Confounding factors that explicitly/implicitly correlate entities. (i) Retrieval: separate causal evidence (behavior, exploit mechanism, toolmarks, infrastructure ownership) from incidental context (sector, region, time), and limit how much context-only metadata can drive retrieval ranking. (ii) Inference: add a counterfactual prompt: “would the conclusion change if we changed sector/region/date?” If yes, lower confidence or abstain, and ask for causal evidence.
1.4 Skewed source. (i) Retrieval: enforce source diversity (e.g., at least $k$ different publishers) and de-duplicate near-identical reports; attach source + timestamp tags to each claim. (ii) Inference: require multi-source support for high-confidence claims; otherwise summarize the disagreement (“A says X, B says Y”) and output a ranked hypothesis list with uncertainty.
1.5 Hierarchical metadata from attack chains. (i) Retrieval: treat kill-chain stages as soft hints unless backed by artifacts (telemetry/logs); attach a confidence note to each stage transition. (ii) Inference: add a strict rule: do not infer later-stage actions unless there is explicit evidence for them; if the chain is partial, keep the answer partial.
Table 7: Ratio of vulnerabilities (same as Figure 2) after applying tailored mitigations. Lower ratio is better.
Method Spurious Correlation Contradictory Knowledge Constrained Generalization
1.1 1.2 1.3 1.4 1.5 2.1 2.2 2.3 2.4 2.5 2.6 3.1 3.2 3.3 3.4
Co-men. Exploit. Confond. Skewed Hier. Temporal Conflict Semantic Diver. Misal. Counter. Distribut. Unseen Overfit. Env.
bias bias factor source metadata contra. report conflict data knowl. CTI bias pattern reasoning unaware
Retrieval-only 3.12 2.05 10.45 2.15 2.11 4.25 6.82 3.55 2.12 1.15 9.45 18.45 19.56 2.55 3.12
General Inference-only 1.85 1.15 5.65 5.23 1.89 10.45 14.56 2.89 3.15 1.95 5.12 12.34 11.45 1.12 5.45
Full-fledged 1.05 0.65 3.12 1.85 0.95 2.55 4.15 1.25 1.55 0.65 3.25 9.56 7.85 0.45 2.15
Retrieval-only 2.56 0.72 13.45 1.12 0.55 2.15 5.12 1.15 1.95 0.45 1.89 19.56 24.12 16.45 6.89
Agent Inference-only 1.25 0.35 7.85 3.45 0.48 6.89 10.45 0.85 3.12 0.65 0.85 13.45 14.56 8.45 11.25
Full-fledged 0.85 0.22 4.56 0.95 0.25 1.55 3.45 0.45 1.25 0.25 0.45 10.56 11.25 5.12 5.45
5.2.2 Mitigating contradictory knowledge failures
Contradictions occur when models encounter conflicting, inconsistent, or incompatible CTI between different threat databases or platforms. These contradictions can arise from differences in temporal scope, analyst interpretation, data formats, or semantic misalignment. To mitigate these failures, systems should incorporate (i) retrieval-based temporal or source-aware annotation and (ii) inference-time conflict detection and uncertainty justification. Detailed strategies are:
2.1 Temporal contradiction between outdated and recent evidence. (i) Retrieval: time-aware reranking (prefer recent + corroborated sources), and package evidence in chronological order with clear as-of time tags. (ii) Inference: when claims conflict, prefer newer evidence unless it is clearly weaker; otherwise present both with timestamps and explain the reason of conflict in a self-reflective generation.
2.2 Conflicting reports of attack contexts or dependencies. (i) Retrieval: cluster retrieved passages into consistent views (by source or by claim similarity) and feed them as grouped bundles rather than a mixed list. (ii) Inference: require a short divergence summary (who claims what) and a decision policy: choose one with justification, or abstain and list what extra evidence would resolve it.
2.3 Semantic conflict. (i) Retrieval: canonicalize entities (aliases, malware family names, tool versions) using an alias table/KB, and attach disambiguation notes when names are similar. (ii) Inference: add a gate before merging claims: “same entity, alias, or related-but-different?” and do not merge unless it is truly the same; also explicitly check for negation (e.g., not attributed).
2.4 Divergent data structure from different platforms. (i) Retrieval: normalize different schemas (STIX, vendor formats) into one simple intermediate record (typed fields or triples) before passing to the model. (ii) Inference: feed the model the normalized record with field definitions, and instruct it to reason from field-specific meaning.
2.5 Misaligned knowledge and security standards. (i) Retrieval: attach authoritative references during retrieval (valid ATT&CK technique IDs, CVSS rubric snippets, NVD-style conventions) and validate candidate mappings early. (ii) Inference: constrain outputs to valid options (pick from a provided list); if evidence does not support a clean mapping, output “unknown/insufficient evidence” with a brief rationale.
2.6 Counteracting CTI generation and LLM alignment. (i) Retrieval: require every major claim to carry a grounding span (quote + source + time), and keep an audit trail of which evidence supported which claim. (ii) Inference: enforce concise bullet claims with a citation to evidence, plus calibrated uncertainty (no confident wording without support).
5.2.3 Mitigating constrained generalization failures
In cyber-security, generalization issues often present as under-performance on emerging threats, low-coverage sectors, or adversarial shifts in attack behavior. To mitigate these issues, systems should incorporate (i) retrieval-based data augmentation from external sources and (ii) inference-time emphasis on generalization rather than memorization:
3.1 Distributional bias. (i) Retrieval: balance exposure by reweighting retrieval results so long-tail actors/regions/tools are not drowned out; add simple quotas or diversity constraints by geography/tooling when possible. (ii) Inference: explicitly warn against common threat actors (or infrastructure) and require the deliberation to cite causal indicators (behavior/exploit/toolmarks), not just region/sector stereotypes.
3.2 Unseen pattern from emerging threats. (i) Retrieval: plug in fresh CTI feeds (recent reports, disclosures, malware analysis notes) with freshness filters and credibility scoring, so the context reflects post-training threats. (ii) Inference: use a novelty-aware response: propose cautious hypotheses, clearly mark what is unknown, and list the top evidence signals that would confirm or refute each hypothesis.
3.3 Overfitted reasoning. (i) Retrieval: diversify evidence so one shortcut (e.g., tool $\rightarrow$ actor) cannot dominate; when possible, retrieve counterevidence or alternatives. (ii) Inference: add an evidence-sufficiency gate: if attribution is not explicitly supported, lower confidence or abstain; optionally run a small counterfactual check that swaps common shortcuts (actor/infra) to see if the conclusion collapses.
3.4 Environmental unawareness. (i) Retrieval: treat environment details as first-class inputs (asset type, deployment, sector constraints, region) and retrieve environment-matched guidance and prior cases. (ii) Inference: require conditional reasoning: “given environment E, risk/mitigation changes because…”, and tie incident response recommendations to environment-specific constraints.
5.2.4 Experiment results on mitigation
The experimental results in Table 7 demonstrate that the above approaches yield consistent reduction in vulnerability rates across all categories, outperforming both ablation baselines (i.e., retrieval-only or inference-only mitigation). A comparative analysis of the ablations reveals distinct component strengths. Retrieval-only mitigation proves superior in resolving contradictory knowledge (Types 2.1, 2.2, 2.4) and environmental unawareness (Type 3.4). This suggests that failures driven by outdated information or source ambiguity are best addressed by grounding the model in retrieved, verifiable evidence (e.g., reducing General-LLM temporal contradiction from 17.53% to 4.25%).
Conversely, Inference-only mitigation dominates in mitigating spurious correlations (Types 1.1, 1.3) and constrained generalization (Types 3.2, 3.3). For instance, in confounding factors (1.3), the inference-based check reduces the failure rate for General LLMs to 5.65% (vs. 10.45% for retrieval), indicating that these errors stem from flawed internal reasoning rather than a lack of information. The full-fledged mitigation further synthesizes retrieval to cleanse the input context and causal inference to filter the reasoning process, thereby achieving the lowest universal failure ratios.
@@ 6 Related Work
LLMs for cybersecurity. LLMs are increasingly leveraged for both offensive and defensive cybersecurity tasks, owing to their strong natural language understanding and reasoning capabilities. In static analysis, LSAST augments traditional SAST tools with dynamic vulnerability knowledge (keltek2025lsast), while hybrid systems use LLM-driven preprocessing and explanation to improve anomaly detection in IoT (ghimire2025enhancing). Autonomous agents have demonstrated the ability to exploit one-day vulnerabilities (fang2024llm) and guide fuzzing across multi-hop library dependencies (zhou2024magneto). Additional applications include enhancing intrusion detection (g2024harnessing), supporting large-scale code review (sun2025bitsai), and enabling malware tracking through dataset augmentation and semantic analysis (yu2024maltracker).
Vulnerabilities of LLMs. Despite their promise, LLM-based agents exhibit critical vulnerabilities that undermine reliability in high-stakes domains. Recent work shows that malfunction amplification attacks can cascade small reasoning errors into severe misjudgments (zhang2024breaking). Other studies highlight vulnerabilities to data poisoning (wang2025comprehensive), adversarial prompts debenedetti2024agentdojo, and misalignment (fang2024preemptive), all of which raise concerns for deploying LLMs in security-sensitive environments.
@@ 7 Conclusion
This work presents intensive studies of intrinsic vulnerabilities that constrain the effectiveness of LLMs in CTI. We conduct large-scale evaluations on CTI benchmark and uncover three fundamental failure modes that persist across multiple CTI stages. We further propose a categorization methodology for in-depth study of failure instances and provides insights into how these limitations emerge and propagate. These findings not only reveal blind spots in current LLM-CTI reasoning but also build a solid initial step to motivate our defensive designs toward robust cybersecurity agents.
Ethical Considerations
Human Subject Research and Annotator Well-being. Our study involves a human-involved categorization framework (Algorithm 1 and 2) to label CTI failure modes. While this includes human efforts, it falls under the category of non-clinical, low-risk technical annotation. All annotators involved in the iterative refinement process and the final verification of failure instances were recruited from professional cybersecurity research teams or were graduate students with domain expertise. The annotated data was public and strictly technical (e.g., TTP descriptions, code snippets, log files) and did not contain offensive, graphic, or sexually explicit content that typically triggers psychological harm concerns.
Data Privacy. The datasets used in this work are derived from two primary sources: (1) Established public benchmarks (CTIBench, SevenLLM-Bench, SWE-Bench, CyberTeam) and (2) Real-world threat reports sourced from public, authoritative repositories (e.g., NVD, CISA advisories, vendor blogs). We ensured that no proprietary, non-public enterprise data was used in the evaluation.
Responsible Disclosure and Dual-Use Concerns. Our research identifies specific vulnerabilities in how LLM-based agents reason about cyber threats (e.g., spurious correlations, contradictory knowledge). While identifying these blind spots could theoretically assist adversaries in crafting adversarial samples to evade automated CTI analysis (e.g., poisoning a threat feed to trigger specific hallucinations), we believe the defensive benefits significantly outweigh the risks. We provide the community with the roadmap to build more resilient agents. Our goal is to highlight architectural limitations in the current generation of models rather than to disclose zero-day vulnerabilities in specific software products. Therefore, standard coordinated disclosure processes were not triggered, as the vulnerabilities are inherent to the probabilistic nature of current LLMs rather than patching flaws in a specific vendor’s code.
Open Science
Availability of Artifacts. To support reproducibility and facilitate future research into resilient CTI agents, we have made artifacts necessary to evaluate our contributions publicly available. We provide the complete implementation of our experimental pipeline. The code is currently hosted at an anonymous repository for double-blind review: https://anonymous.4open.science/r/LLM-CTI-E773. Upon acceptance of this paper, the code will be migrated to a permanent, public GitHub repository with a DOI via Zenodo, as well as MIT License to encourage broad adoption and modification.
@@ Appendix A Complementary Details of CTI
  • Topic Modeling. Topic modeling techniques such as Latent Dirichlet Allocation (LDA) [blei2003lda], Non-negative Matrix Factorization [lee1999nmf], and more recent neural topic models [dieng2020topic] are applied to large corpora of unstructured CTI text, including threat reports, incident tickets, and log annotations. These models group documents or paragraphs into coherent themes, enabling analysts to identify clusters of related activity such as “phishing campaigns leveraging Office macros” or “ransomware families exploiting VPN vulnerabilities.” In practice, topic modeling supports early triage by prioritizing threat feeds that share thematic overlap with active campaigns in a given sector.
  • Event Extraction. Event extraction identifies structured incidents—who did what, when, and how—from unstructured logs or reports. Natural Language Processing (NLP) pipelines detect entities such as vulnerabilities (CVEs), indicators of compromise (IP addresses, file hashes), and TTPs (MITRE ATT&CK techniques [strom2018mitre]), then associate them with temporal markers and actor actions. In Security Operations Centers (SOCs), event extraction enables long-form reports to be converted into structured JSON objects or STIX bundles [oasis2017stix], which can be automatically ingested into SIEM platforms such as Splunk or Elastic. This reduces manual parsing effort and ensures consistent representation across heterogeneous data sources.
  • Knowledge Base Mapping. Knowledge base mapping aligns extracted entities and events with standardized taxonomies such as CVE [cve], CWE [cwe], MITRE ATT&CK [strom2018mitre], and CAPEC [capec]. This process typically involves both exact matching (e.g., direct CVE ID resolution) and approximate entity linking (e.g., mapping the phrase “remote PowerShell execution” to ATT&CK T1059.001). In CTI practice, knowledge base mapping is essential for interoperability: intelligence can be shared across organizations using a common language of identifiers, enabling correlation of local incidents with global adversary behaviors.
  • Information Retrieval. Information retrieval systems allow analysts to ground their findings in relevant historical reports and external databases. Implementations include keyword- and embedding-based retrieval from proprietary CTI feeds (e.g., Recorded Future, Mandiant Advantage), open-source repositories (e.g., AlienVault OTX, AbuseIPDB), and structured vulnerability catalogs such as NVD [nvd] and the CISA KEV catalog [CISA_KEV_Catalog]. Retrieval pipelines often combine lightweight keyword filters for precision with dense embedding search for semantic coverage. Analysts use these systems to verify whether a newly observed domain has prior associations with known malware campaigns, or whether a vulnerability is actively being exploited.
@@
  • Named Entity Recognition (NER). NER systems are used to extract structured entities such as malware families, infrastructure elements, or actor names from unstructured reports and incident logs. For example, extracting references to APT28, Mimikatz, or QakBot across heterogeneous feeds helps analysts consolidate threat narratives [barnes2021cyberner].
  • Relation and Event Extraction. Beyond isolated entities, attribution requires uncovering relationships among adversary tactics, techniques, and infrastructure. Relation extraction techniques map, for instance, shared IP ranges between phishing campaigns or code reuse across malware variants, while event extraction captures sequences of actions such as exploitation followed by lateral movement [husari2018tactic].
  • Infrastructure Correlation. Adversaries frequently reuse or repurpose command-and-control (C2) infrastructure. CTI teams apply graph-based correlation to link domain registrations, TLS certificates, and hosting providers, enabling attribution of otherwise fragmented observations to known actor toolkits or campaigns [stonegross2009your, antonakakis2017understanding].
  • Stylistic and Linguistic Profiling. Analysts also consider linguistic cues in adversary communications or malware code artifacts. Stylometry and compilation fingerprinting methods can identify patterns in variable naming, debugging strings, or grammar usage, which help to tie malware development back to particular groups [rosenblum2011learning].
  • Campaign Graph Construction. To synthesize diverse evidence, CTI analysts build structured graphs of adversarial campaigns, linking entities, infrastructure, and TTPs across incidents. Graph construction enables propagation of attribution hypotheses and detection of actor evolution over time [yang2021graphbased].
@@
  • Historical correlation and trend analysis. Analysts correlate prior incidents and intrusion campaigns to identify recurring attacker playbooks. For example, statistical methods on longitudinal CVE exploitation data help assess whether recently disclosed vulnerabilities follow exploitation trends of past families [Bilge2012ZeroDay, Allodi2014Economic].
  • Exploit prediction scoring. Models such as the Exploit Prediction Scoring System (EPSS) estimate the probability that a vulnerability will be exploited within a given time window, using features such as vulnerability metadata, CVSS scores, and real-world exploit observations [Jacobs2019EPSSarxiv, Jacobs2021EPSS].
  • Temporal modeling of campaign progression. Recurrent neural networks and temporal point processes capture how campaigns unfold over time, modeling likely transitions from initial access to follow-on payloads such as ransomware [Okutan2020TemporalThreat].
  • Temporal forcasting. LLMs predict which TTPs an actor is likely to employ next or what’s impact. This supports proactive defense, such as generating detection rules for tactics not yet observed in the ongoing campaign [Sood2013APT, Husari2018ThreatModeling].
  • Threat propagation simulation (synthesis). Agent-based and epidemic-style models simulate the spread of malware or worms across interconnected systems, forecasting infection curves and propagation likelihood [Liu2015MalwarePropagation].
@@
  • Mitigation efficacy evaluation. Beyond applying countermeasures, analysts must assess their effectiveness. Approaches include red-teaming exercises, breach-and-attack simulation (BAS) platforms, and adversary emulation scenarios that replay known TTPs to test whether mitigations succeed in preventing, detecting, or containing malicious activity [shostack2014threat, peisert2021towards].
  • Response playbook recommendation. Structured playbooks standardize incident response by encoding lessons from CTI. These include step-by-step containment and recovery actions tailored to adversary campaigns (e.g., disabling compromised accounts, isolating infected subnets). Orchestration tools such as SOAR platforms automate playbook execution, integrating CTI feeds into dynamic workflows [mitnick2018art, scarfone2012computer].
  • Summarization. Finally, mitigation intelligence is communicated through concise, contextualized reports for executives and IT operators. Summarization synthesizes prioritized vulnerabilities, mapped mitigations, and recommended workflows. This ensures decision makers understand trade-offs between operational impact and security gains, and supports cross-team coordination in enterprise-scale defense [mavroeidis2018cyber].